OpenStudy (anonymous):

For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? -3/2 -2/3 2 3

OpenStudy (jackellyn):

To find the x intercept, set y = 0 and solve for x. 2x^2 - 7x + 6 =0 (x-2)(2x-3)=0 If you set each of these equal to 0 and solve for x then it will give you the intercepts. Can you do that?

OpenStudy (anonymous):

ya thanks

OpenStudy (anonymous):

set the x to 0?

OpenStudy (anonymous):

you there?

OpenStudy (jackellyn):

No set it equal to 0. x-2=0 , 2x-3=0 And solve for x on both